# Improved Scaling Laws in Linear Regression via Data Reuse

### v1

@@ B.3 Bias error under the source condition
Let Assumption 1 hold, $a>b-1$, and assume ${L_{\mathtt{eff}}}\lesssim N^{a}/\gamma$. Under the notation in Theorem 3.1 and its proof, there exist some $(a,b)$-dependent constants $c,c^{\prime}>0$ such that when $\gamma\leq c/\log N$,
@@
[…]
@@
[…]
@@ C.1 Upper and lower bounds
[…]
@@
and $(\widehat{\lambda}_{i})_{i=1}^{M}$ are the eigenvalues of ${\bm{\widehat{\Sigma}}}$.
@@
Adopt the shorthand $\mathsf{V}_{L}$ for $\mathbf{I}-\prod_{t=1}^{L}(\mathbf{I}-\gamma_{t}{\bm{\widehat{\Sigma}}})$. Reorganizing the terms, we have
@@
Let $\widehat{\lambda}_{1},\ldots,\widehat{\lambda}_{M}$ be the eigenvalues of ${\bm{\widehat{\Sigma}}}$ in non-increasing order, and let $\lambda>0$ be some value which will be given later. We now derive an upper bound and a lower bound for the variance $\mathsf{Var}.$
@@
Similar to the proof of claim (7b) in Lemma B.2, it can be verified that $\mathsf{V}_{L}\preceq\mathbf{I}-(\mathbf{I}-2\gamma_{0}{\bm{\widehat{\Sigma}}})^{L_{\mathtt{eff}}}$ under the condition $\gamma_{0}\leq 1/[4\max_{i}\|\mathbf{S}\mathbf{x}_{i}\|_{2}^{2}]\leq 1/[4\operatorname{tr}({\bm{\widehat{\Sigma}}})]$ and stepsize assumption (2). Since $(1-(1-\gamma_{0}x)^{L_{\mathtt{eff}}})^{2}\leq\min\{(x{L_{\mathtt{eff}}}\gamma_{0})^{2},1\}$ for $x\in[0,1/(2\gamma_{0})]$ by Bernoulli’s inequality and $\sup_{[0,1/\gamma_{0}]}[(1-(1-\gamma_{0}x)^{L_{\mathtt{eff}}})/x]={L_{\mathtt{eff}}}\gamma_{0},$ it follows that
@@
#$\lesssim\#\{i\in[M]:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}>1/4\}+({L_{\mathtt{eff}}}\gamma_{0})^{2}\sum_{i:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}\leq 1/4}\widehat{\lambda}_{i}^{2}+({L_{\mathtt{eff}}}\gamma_{0})\sum_{i:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}\leq 1/4}\widehat{\lambda}_{i}=:\widetilde{D}^{\mathsf{U}}$
#$\lesssim\#\{i\in[M]:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}>1/4\}+({L_{\mathtt{eff}}}\gamma_{0})\sum_{i:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}\leq 1/4}\widehat{\lambda}_{i}=:\widetilde{D}^{\mathsf{U}}$ (16)
Applying Lemma E.1 and noting that $\operatorname{tr}(\mathsf{V}_{L}^{2}+\lambda{\bm{\widehat{\Sigma}}}^{-1}\mathsf{V}_{L}^{2})\lesssim N$ as $\mathsf{V}_{L}$ has at most $N$ non-zero eigenvalues, we obtain
@@
=$\overset{(ii)}{=}\sum_{i=1}^{M}\mu_{i}({\bm{\Sigma}})\cdot\mu_{i}({\bm{\widehat{\Sigma}}})\cdot\mu_{2(M-i)+1}(\mathsf{V}_{L}^{2}{\bm{\widehat{\Sigma}}}^{-2}),$
@@
Note that $\mathsf{V}_{L}\succeq\mathbf{I}-(\mathbf{I}-\gamma_{0}{\bm{\widehat{\Sigma}}})^{L_{\mathtt{eff}}}$. Since $f(x):=(1-(1-\gamma_{0}x)^{{L_{\mathtt{eff}}}})^{2}/x^{2}$ is a decreasing function on $[0,1/\gamma_{0}]$ and (1) $f(x)\geq({L_{\mathtt{eff}}}\gamma_{0})^{2}/4$ when ${L_{\mathtt{eff}}}\gamma_{0}x\leq 1/4$; (2) $f(x)\geq 1/(5x^{2})$ when ${L_{\mathtt{eff}}}\gamma_{0}x\geq 1/4$, we have
@@ C.2 Variance error under the source condition
$D^{\mathsf{L}}$ $\geq{\mathbb{E}}_{\mathbf{X}}\Bigg{[}({L_{\mathtt{eff}}}\gamma_{0})^{2}\cdot\sum_{i:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}\leq 1/4}\mu_{i}({\bm{\Sigma}})\cdot\mu_{i}({\bm{\widehat{\Sigma}}})\Bigg{]}$
@@ D.1 An upper bound
$\bm{\theta}^{{(-i)}}_{t}=(\mathbf{I}-\gamma_{t}{\bm{\widehat{\Sigma}}}^{{(-i)}})\bm{\theta}^{{(-i)}}_{t-1}+\gamma_{t}(\mathbf{S}\mathbf{X}^{\top}\mathbf{y})^{{(-i)}},~{}~{}\text{with~{}~{}}\bm{\theta}^{{(-i)}}_{0}=\bm{0},$ (LOO-GD)
@@
Moreover, if $\mu_{j}({\bm{\widehat{\Sigma}}})\eqsim j^{-a}$ for $j\leq r({\bm{\widehat{\Sigma}}})$ for some $a>1$, then
@@
Define $\bm{\Delta}_{t}:=\mathbf{v}_{t}-\bm{\theta}_{t}$ for $t\in[0,L]$. Recall that $\mathbf{v}^{*}=(\mathbf{S}\mathbf{H}\mathbf{S}^{\top})^{-1}\mathbf{S}\mathbf{H}\mathbf{w}^{*}$ and we have $y_{i}=(\mathbf{S}\mathbf{x}_{i})^{\top}\mathbf{v}^{*}+\tilde{\epsilon}_{i}$ for all $i\in[N]$ with $\tilde{\epsilon}_{i}$ independent of $\mathbf{S}\mathbf{x}_{i}$ conditioned on $(\mathbf{S},\mathbf{w}^{*})$ under the Gaussian assumption in Assumption 1A. Moreover, ${\mathbb{E}}[\tilde{\epsilon}_{i}|\mathbf{S},\mathbf{w}^{*}]=0$ and ${\mathbb{E}}[\tilde{\epsilon}_{i}^{2}|\mathbf{S},\mathbf{w}^{*}]\leq\sigma^{2}+\|\mathbf{w}^{*}\|^{2}_{\mathbf{H}}=:\tilde{\sigma}^{2}(\mathbf{w}^{*})$. By the definition of $\mathbf{v}_{t}$ and $\bm{\theta}_{t}$ in (multi-pass SGD) and (GD), we have
@@
Note that conditioned on $\mathbf{w}^{*}$, $\mathbf{S}$ and the dataset ${\mathcal{D}}=(\mathbf{x}_{i},y_{i})_{i=1}^{N}$, the noise terms ${\mathbb{E}}[\bm{\xi}_{1,t}|\mathbf{S},\mathbf{w}^{*},{\mathcal{D}}]={\mathbb{E}}[\bm{\xi}_{2,t}|\mathbf{S},\mathbf{w}^{*},{\mathcal{D}}]=0$. The proof of this Lemma follows from similar ideas as in the proof of Lemma 5 in Pillaud-Vivien et al. (2018), but with a more precise characterization on the magnitude of GD outputs. We start with presenting the following two results.
@@
${\mathbb{E}}_{i_{t},t\in[L]}\|{\bm{\Sigma}}^{1/2}(\mathbf{v}_{L}-\bm{\theta}_{L})\|_{2}^{2}$ $\lesssim\|({\bm{\widehat{\Sigma}}}+\lambda\mathbf{I})^{-1/2}({\bm{\Sigma}}+\lambda\mathbf{I})^{1/2}\|^{2}\cdot{\mathbb{E}}_{i_{t},t\in[L]}\|({\bm{\widehat{\Sigma}}}+\lambda\mathbf{I})^{1/2}(\mathbf{v}_{L}-\bm{\theta}_{L})\|_{2}^{2}$
@@
Conditioned on $\mathbf{S}$ and $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$, when choosing $\bm{\nu}_{t}=\mathbf{S}\mathbf{x}_{i_{t}}$, we have ${\mathbb{E}}[\bm{\nu}_{t}\bm{\nu}_{t}^{\top}]={\mathbb{E}}[\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top}]={\bm{\widehat{\Sigma}}}$, and ${\mathbb{E}}[\bm{\nu}_{t}\bm{\nu}_{t}^{\top}\bm{\nu}_{t}\bm{\nu}_{t}^{\top}]\leq{\mathbb{E}}[\max_{i\in[N]}\|\bm{\nu}_{i}\|^{2}\bm{\nu}_{i}\bm{\nu}_{i}^{\top}]\preceq\max_{i\in[N]}\|\mathbf{S}\mathbf{x}_{i}\|_{2}^{2}\cdot{\bm{\widehat{\Sigma}}}=B_{\bm{\nu}}^{2}\cdot{\bm{\widehat{\Sigma}}}$. Thus we may let ${\bm{\Sigma}}_{\bm{\nu}}={\bm{\widehat{\Sigma}}}$ and $B_{\bm{\nu}}=\max_{i\in[N]}\|\mathbf{S}\mathbf{x}_{i}\|_{2}$. In this case, we have $\gamma_{0}B_{\bm{\nu}}^{2}\leq 1/4$ by the assumption that $\gamma_{0}\leq 1/[4\max_{i\in[N]}\|\mathbf{S}\mathbf{x}_{i}\|_{2}^{2}]$. It remains to bound $\sigma_{\xi}^{2}$ in (19). Note that
@@ D.2 A lower bound
where $B_{\bm{\xi}}:=\max\{\frac{N-1}{N}-\frac{4\gamma_{0}{L_{\mathtt{eff}}}}{N}B_{\bm{\nu}}^{2},0\}$, $B_{\bm{\nu}}:=\max_{i\in[N]}\|\mathbf{S}\mathbf{x}_{i}\|_{2}$ and $\tilde{t}:=\#\{i\in[M]:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}>1/8\}$, and $(\widehat{\lambda}_{i})_{i=1}^{M}$ are the eigenvalues of ${\bm{\widehat{\Sigma}}}$.
@@
$\succeq\tilde{\sigma}^{2}(\mathbf{w}^{*})B_{\bm{\xi}}\cdot\sum_{i=1}^{L}\gamma_{i}^{2}{\mathbb{E}}_{i_{t},t\in[L]}[\operatorname{tr}(\prod_{j=i+1}^{L}(\mathbf{I}-\gamma_{j}\mathbf{S}\mathbf{x}_{i_{j}}\mathbf{x}_{i_{j}}^{\top}\mathbf{S}^{\top}){\bm{\Sigma}}\prod_{j=i+1}^{L}(\mathbf{I}-\gamma_{j}\mathbf{S}\mathbf{x}_{i_{j}}\mathbf{x}_{i_{j}}^{\top}\mathbf{S}^{\top})^{\top}{\bm{\widehat{\Sigma}}})]$
$\succeq\tilde{\sigma}^{2}(\mathbf{w}^{*})B_{\bm{\xi}}\cdot\sum_{i=1}^{L}\gamma_{i}^{2}\operatorname{tr}({\bm{\Sigma}}{\bm{\widehat{\Sigma}}}\prod_{j=i+1}^{L}(\mathbf{I}-2\gamma_{j}{\bm{\widehat{\Sigma}}})),$
where the last line follows from the fact that ${\mathbb{E}}_{i_{j}}[(\mathbf{I}-\gamma_{j}\mathbf{S}\mathbf{x}_{i_{j}}\mathbf{x}_{i_{j}}^{\top}\mathbf{S}^{\top})P({\bm{\widehat{\Sigma}}})(\mathbf{I}-\gamma_{j}\mathbf{S}\mathbf{x}_{i_{j}}\mathbf{x}_{i_{j}}^{\top}\mathbf{S}^{\top})]\succeq P({\bm{\widehat{\Sigma}}})(\mathbf{I}-2\gamma_{j}{\bm{\widehat{\Sigma}}})$ for any polynomial $P$. Continuing from the last line, we have
@@
$=\mu_{M-j}({\bm{\widehat{\Sigma}}})\cdot\mu_{M-j}({\bm{\Sigma}})\cdot\mu_{3j-2}\Big{(}\frac{(\mathbf{I}-(\mathbf{I}-2\gamma_{0}{\bm{\widehat{\Sigma}}})^{{L_{\mathtt{eff}}}})(\mathbf{I}-4\gamma_{0}{\bm{\widehat{\Sigma}}})^{{L_{\mathtt{eff}}}}}{{\bm{\widehat{\Sigma}}}}\Big{)}.$
@@
Since ${\mathbb{E}}_{\tilde{\bm{\epsilon}}}[(\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top}-{\bm{\widehat{\Sigma}}})\prod_{i=1}^{t}(\mathbf{I}-\gamma_{i}{\bm{\widehat{\Sigma}}})\mathbf{v}^{*}\tilde{\bm{\epsilon}}^{\top}]=\bm{0}$, it can be verified that
@@
$\succeq\tilde{\sigma}^{2}(\mathbf{w}^{*})\cdot\Big{[}\frac{N-1}{N}{\bm{\widehat{\Sigma}}}-(\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top}-{\bm{\widehat{\Sigma}}})\Big{[}\frac{\mathbf{I}-\prod_{i=1}^{t}(\mathbf{I}-\gamma_{i}{\bm{\widehat{\Sigma}}})^{2}}{N{\bm{\widehat{\Sigma}}}}\Big{]}(\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top}-{\bm{\widehat{\Sigma}}})^{\top}\Big{]}$
$\succeq\tilde{\sigma}^{2}(\mathbf{w}^{*})\cdot\Big{[}\frac{N-1}{N}{\bm{\widehat{\Sigma}}}-(\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top}-{\bm{\widehat{\Sigma}}})\Big{[}\frac{\mathbf{I}-(\mathbf{I}-2\gamma_{0}{\bm{\widehat{\Sigma}}})^{2{L_{\mathtt{eff}}}}}{N{\bm{\widehat{\Sigma}}}}\Big{]}(\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top}-{\bm{\widehat{\Sigma}}})^{\top}\Big{]}.$
Since $\sup_{x\in[0,1/(2\tilde{\gamma})]}(1-(1-2\tilde{\gamma}x)^{2{L_{\mathtt{eff}}}})/x\leq 4\tilde{\gamma}{L_{\mathtt{eff}}}$ and ${\mathbb{E}}_{i_{t}}[(\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top}-{\bm{\widehat{\Sigma}}})^{2}]\preceq{\mathbb{E}}_{i_{t}}[(\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top})^{2}]\preceq B_{\bm{\nu}}^{2}{\bm{\widehat{\Sigma}}}$, we further have

### v2

@@ B.3 Bias error under the source condition
Let Assumption 1 hold, a>b1a>b-1, and assume ${L_{\mathtt{eff}}}\lesssim N^{a}/\gamma$. Under the notation in Theorem 3.1 and its proof, there exist some $(a,b)$-dependent constants $c,c^{\prime}>0$ such that when $\gamma\leq c/\log N$,
@@
To verify condition (4) in Assumption 3, we introduce $\mathbf{z}_{1},\ldots\mathbf{z}_{N}\overset{iid}{\sim}\mathcal{N}(0,\mathbf{I}_{M}/N)$ and write $\mathbf{Z}=(\mathbf{z}_{1},\ldots,\mathbf{z}_{N})^{\top}$. It can be shown that $\mathbf{X}\mathbf{S}^{\top}/\sqrt{N}\overset{d}{=}\mathbf{Z}{\bm{\Sigma}}^{1/2}$ conditioned on $\mathbf{S}$. Therefore, we have $\operatorname{tr}({\bm{\widehat{\Sigma}}})\overset{d}{=}\operatorname{tr}(\mathbf{Z}{\bm{\Sigma}}\mathbf{Z}^{\top}),$ where ${\mathbf{Z}}\in{\mathbb{R}}^{N\times M}$ is a Gaussian sketching matrix. When $\mu_{i}({\bm{\Sigma}})\eqsim i^{-a}$ for all $i\in[M]$ (which happens with probability at least $1-\exp(-\Omega(M))$ over $\mathbf{S}$ by Lemma E.5), we have by Hansen-Wright inequality (see e.g., exercise 2.17 in Wainwright (2019)) and a union bound that
@@
with probability at least $1-e^{-\Omega(M)}$, where $\overline{t}={\mathbb{E}}_{\mathbf{X}}[\#\{i\in[M]:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}>1/4\}]$ and $\{\widehat{\lambda}_{i},i\in[M]\}$ are the eigenvalues of $\bm{\widehat{\Sigma}}$}. Since ${\bm{\widehat{\Sigma}}}\overset{d}{=}\mathbf{Z}{\bm{\Sigma}}\mathbf{Z}^{\top}$ conditioned on $\mathbf{S}$ by Lemma E.3, when $\mu_{i}({\bm{\Sigma}})\eqsim i^{-a}$ for all $i\in[M]$ (which happens with probability at least $1-\exp(-\Omega(M))$ over $\mathbf{S}$ by Lemma E.5), we have by combining Lemma E.4 and E.3 with $k=N/c$ that
$\widehat{\lambda}_{2j-1}$ $\overset{d}{=}\mu_{2j-1}(\mathbf{Z}{\bm{\Sigma}}\mathbf{Z}^{\top})$
@@ C.1 Upper and lower bounds
$\mathsf{Var}$ $:={\mathbb{E}}[\operatorname{tr}(\mathbf{X}\mathbf{S}^{\top}\mathbf{V}({\bm{\widehat{\Sigma}}}){\bm{\Sigma}}\mathbf{V}({\bm{\widehat{\Sigma}}})\mathbf{S}\mathbf{X}^{\top})]\lesssim\frac{D^{\mathsf{U}}}{N},~\text{~~and~~~}\mathsf{Var}\gtrsim\frac{D^{\mathsf{L}}}{N},$
@@
and $(\widehat{\lambda}_{i})_{i=1}^{M}$ are the eigenvalues of $^{\bm{\widehat{\Sigma}}}$.
@@
[…]
@@
Let $\widehat{\lambda}_{1},\ldots,\widehat{\lambda}_{M}$ be the eigenvalues of $^{\bm{\widehat{\Sigma}}}$ in non-increasing order, and let $\lambda>0$ be some value which will be given later. We now derive an upper bound and a lower bound for the variance $\mathsf{Var}.$
@@
[…]
@@
$\lesssim\#\{i\in[M]:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}>1/4\}+({L_{\mathtt{eff}}}\gamma_{0})^{2}\sum_{i:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}\leq 1/4}\widehat{\lambda}_{i}^{2}+({L_{\mathtt{eff}}}\gamma_{0})\sum_{i:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}\leq 1/4}\widehat{\lambda}_{i}=:\widetilde{D}^{\mathsf{U}}$
$\lesssim\#\{i\in[M]:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}>1/4\}+({L_{\mathtt{eff}}}\gamma_{0})\sum_{i:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}\leq 1/4}\widehat{\lambda}_{i}=:\widetilde{D}^{\mathsf{U}}$ (17)
Applying Lemma E.1 and noting that $\operatorname{tr}(\mathsf{V}_{L}^{2}+\lambda{\bm{\widehat{\Sigma}}}^{-1}\mathsf{V}_{L}^{2})\lesssim N$ as $\mathsf{V}_{L}$ has at most NN non-zero eigenvalues, we obtain
@@
$\overset{(ii)}{=}\sum_{i=1}^{M}\mu_{i}({\bm{\Sigma}})\cdot\mu_{i}({\bm{\widehat{\Sigma}}})\cdot\mu_{2(M-i)+1}(\mathsf{V}_{L}^{2}{\bm{\widehat{\Sigma}}}^{-2}),$
@@
Note that $\mathsf{V}_{L}\succeq\mathbf{I}-(\mathbf{I}-\gamma_{0}{\bm{\widehat{\Sigma}}})^{L_{\mathtt{eff}}}$. Since $f(x):=(1-(1-\gamma_{0}x)^{{L_{\mathtt{eff}}}})^{2}/x^{2}$ is a decreasing function on $0,1/\gamma_{0}]$ and (1) $f(x)\geq({L_{\mathtt{eff}}}\gamma_{0})^{2}/4$ when ${L_{\mathtt{eff}}}\gamma_{0}x\leq 1/4$; (2) $f(x)\geq 1/(5x^{2})$ when ${L_{\mathtt{eff}}}\gamma_{0}x\geq 1/4$, we have
@@ C.2 Variance error under the source condition
[…]
@@ D.1 An upper bound
[…]
@@
Moreover, if $\mu_{j}({\bm{\widehat{\Sigma}}})\eqsim j^{-a}$ $j\leq r({\bm{\widehat{\Sigma}}}) for$ some $a>1$, then
@@
The proof of this lemma follows from similar ideas as in the proof of Lemma 5 in Pillaud-Vivien et al. (2018), but with a more precise characterization on the magnitude of GD outputs. We start with an overview of the proof. At a high level, to bound the fluctuation error, we express the difference between the multi-pass SGD and GD trajectories, $\mathbf{v}_{t}-{\bm{\theta}}_{t}$, as a stochastic process (Eq. 18) that fits into the framework of Lemma D.2, which provides an upper bound on the fluctuation error $\mathbb{E}[\|{{\bm{\Sigma}}}^{1/2}(\mathbf{v}_{t}-\bm{\theta}_{t})\|^{2}]$ under certain conditions, up to a mismatch between $\widehat{{{\bm{\Sigma}}}}$ and $\bm{\Sigma}$}}. We verify that the required conditions hold with appropriate choices of parameters (Eq. 20), which are further bounded using a leave-one-out argument (Lemma D.3). Applying Lemma D.2 with these parameters and a covariance replacement trick (Eq. 21) yields the desired bounds.
We now proceed to the proof. Define $\bm{\Delta}_{t}:=\mathbf{v}_{t}-\bm{\theta}_{t}$ for $t\in[0,L]$. Recall that $\mathbf{v}^{*}=(\mathbf{S}\mathbf{H}\mathbf{S}^{\top})^{-1}\mathbf{S}\mathbf{H}\mathbf{w}^{*}$ and we have $y_{i}=(\mathbf{S}\mathbf{x}_{i})^{\top}\mathbf{v}^{*}+\tilde{\epsilon}_{i}$ for all $i\in[N]$ with $\tilde{\epsilon}_{i}$ independent of $\mathbf{S}\mathbf{x}_{i}$ conditioned on $(\mathbf{S},\mathbf{w}^{*})$ under the Gaussian assumption in Assumption 1A. Moreover, ${\mathbb{E}}[\tilde{\epsilon}_{i}|\mathbf{S},\mathbf{w}^{*}]=0$ and ${\mathbb{E}}[\tilde{\epsilon}_{i}^{2}|\mathbf{S},\mathbf{w}^{*}]\leq\sigma^{2}+\|\mathbf{w}^{*}\|^{2}_{\mathbf{H}}=:\tilde{\sigma}^{2}(\mathbf{w}^{*})$. By the definition of $\mathbf{v}_{t}$ and $\bm{\theta}_{t}$ in (multi-pass SGD) and (GD), we have
@@
Note that conditioned on $\mathbf{w}^{*}$, $\mathbf{S}$ and the dataset ${\mathcal{D}}=(\mathbf{x}_{i},y_{i})_{i=1}^{N}$, the noise terms ${\mathbb{E}}[\bm{\xi}_{1,t}|\mathbf{S},\mathbf{w}^{*},{\mathcal{D}}]={\mathbb{E}}[\bm{\xi}_{2,t}|\mathbf{S},\mathbf{w}^{*},{\mathcal{D}}]=0$. Next, we present the following two results.
@@
${\mathbb{E}}_{i_{t},t\in[L]}\|{\bm{\Sigma}}^{1/2}(\mathbf{v}_{L}-\bm{\theta}_{L})\|_{2}^{2}$ $\lesssim\|({\bm{\widehat{\Sigma}}}+\lambda\mathbf{I})^{-1/2}({\bm{\Sigma}}+\lambda\mathbf{I})^{1/2}\|^{2}\cdot{\mathbb{E}}_{i_{t},t\in[L]}\|({\bm{\widehat{\Sigma}}}+\lambda\mathbf{I})^{1/2}(\mathbf{v}_{L}-\bm{\theta}_{L})\|_{2}^{2}$ (21)
@@
Conditioned on $\mathbf{S}$ and $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$, when choosing $\bm{\nu}_{t}=\mathbf{S}\mathbf{x}_{i_{t}}$, we have $^{\mathbb{E}}[\bm{\nu}_{t}\bm{\nu}_{t}^{\top}]={\mathbb{E}}[\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top}]={\bm{\widehat{\Sigma}}}$, and $^{\mathbb{E}}[\bm{\nu}_{t}\bm{\nu}_{t}^{\top}\bm{\nu}_{t}\bm{\nu}_{t}^{\top}]\leq{\mathbb{E}}[\max_{i\in[N]}\|\bm{\nu}_{i}\|^{2}\bm{\nu}_{i}\bm{\nu}_{i}^{\top}]\preceq\max_{i\in[N]}\|\mathbf{S}\mathbf{x}_{i}\|_{2}^{2}\cdot{\bm{\widehat{\Sigma}}}=B_{\bm{\nu}}^{2}\cdot{\bm{\widehat{\Sigma}}}$. Thus we may let $^{\bm{\Sigma}}_{\bm{\nu}}={\bm{\widehat{\Sigma}}}$ and $B_{\bm{\nu}}=\max_{i\in[N]}\|\mathbf{S}\mathbf{x}_{i}\|_{2}$. In this case, we have $\gamma_{0}B_{\bm{\nu}}^{2}\leq 1/4$ by the assumption that $\gamma_{0}\leq 1/[4\max_{i\in[N]}\|\mathbf{S}\mathbf{x}_{i}\|_{2}^{2}]$. It remains to bound $\sigma_{\xi}^{2}$ in (20). Note that
@@ D.2 A lower bound
where $B_{\bm{\xi}}:=\max\{\frac{N-1}{N}-\frac{4\gamma_{0}{L_{\mathtt{eff}}}}{N}B_{\bm{\nu}}^{2},0\}$, $B_{\bm{\nu}}:=\max_{i\in[N]}\|\mathbf{S}\mathbf{x}_{i}\|_{2}$ and $\tilde{t}:=\#\{i\in[M]:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}>1/8\}$, and $(\widehat{\lambda}_{i})_{i=1}^{M}$ are the eigenvalues of $^{\bm{\widehat{\Sigma}}}$.
@@
[…]
$),\displaystyle\succeq\tilde{\sigma}^{2}(\mathbf{w}^{*})B_{\bm{\xi}}\cdot\sum_{i=1}^{L}\gamma_{i}^{2}\operatorname{tr}({\bm{\Sigma}}{\bm{\widehat{\Sigma}}}\prod_{j=i+1}^{L}(\mathbf{I}-2\gamma_{j}{\bm{\widehat{\Sigma}}}$)),
[…]
@@
[…]
@@
Since ${\mathbb{E}}_{i_{t},t\in[L]}[\bm{\xi}^{s}_{i}]=0$, it can be verified that
@@
[…]
Since $\sup_{x\in[0,1/(2\tilde{\gamma})]}(1-(1-2\tilde{\gamma}x)^{2{L_{\mathtt{eff}}}})/x\leq 4\tilde{\gamma}{L_{\mathtt{eff}}}$ and $^{\mathbb{E}}_{i_{t}}[(\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top}-{\bm{\widehat{\Sigma}}})^{2}]\preceq{\mathbb{E}}_{i_{t}}[(\mathbf{S}\mathbf{x}_{i_{t}}\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top})^{2}]\preceq B_{\bm{\nu}}^{2}{\bm{\widehat{\Sigma}}}$, we further have